# Mixture of Group Experts for Learning Invariant Representations

### v1

@@ 4.2. Language Modeling
Setup. Following (Teo and Nguyen, 2024), we evaluate the effectiveness of MoGE on WikiText-103 (Merity et al., 2017) using three types of MoE models: Switch Transformer (SMoE) (Fedus et al., 2022), MomentumSMoE (Teo and Nguyen, 2024), and AdamSMoE (Teo and Nguyen, 2024). Each MoE layer consists of 16 experts utilizing top-2 routing. MoGE introduces our proposed group sparse regularization to the MoE, while keeping all other aspects unchanged.
@@
Results. Table 3 summarizes the perplexity (PPL) values for three types of MoE models on the validation and the test sets of WikiText-103. Lower PPL values indicate better performance, with the best results highlighted in bold. MoGE consistently outperforms MoE across all test sets and most validation sets, clearly demonstrating its superiority. Although MoGE shows slightly inferior performance to MoE on the validation set when using the AdamSMoE-medium model, it still achieves better performance on the more critical test set. Due to space constraints, the training and the validation PPL curves are provided in Appendix A.
Table 3. Comparison of the perplexity (PPL) of MoGE across three types of MoE models. The down-arrow notation ($\downarrow$) indicates that a lower value represents better performance.
@@
4.3. Scalability and Efficiency Analysis
Scaling $k$ and $n$. We analyze the performance of MoE and MoGE when scaling the number of experts (i.e., $n$) and the $k$ value in top-$k$ routing. For this investigation, we use the V-MoE-S/8 and V-MoGE-S/8 models, trained from scratch on Tiny-ImageNet, as in Table 2. The results, summarized in Table 4, consistently demonstrate that MoGE outperforms MoE. Moreover, MoGE exhibits a less significant performance decline as $n$ and $k$ increase. This highlights the capacity of MoGE to leverage a higher number of experts and larger $k$ values, providing substantial evidence for the flexible and scalable design of its experts.
Additional Computation and Memory. Compared to MoE, the additional computational and memory overhead introduced by MoGE primarily depends on the number of experts (i.e., $n$), rather than the size of the experts or the total parameter count of the MoE model. To empirically validate this fact, we select four models from Table 2 and evaluate them under two scenarios: training from scratch and fine-tuning, with $n=32$ and top-1 routing. The results are summarized in Table 5.
The increase in total GPU memory usage for MoGE, compared to MoE, is negligible. When comparing the wall-clock time, we observe that for training ViT-based MoE models from scratch on Tiny-ImageNet, the training time for MoGE increases by no more than 1% compared to MoE. Similarly, for fine-tuning SwinV1-based MoE models on ImageNet-1K, the training time for MoGE increases by no more than 1.5% compared to MoE. Although ViT-S/8 and ViT-B/8 differ significantly in total model parameters, they share the same number of MoE layers, resulting in a comparable absolute increase in training time (248 vs. 254 seconds). Likewise, SwinV-T/7 and SwinV1-B/7 differ in parameter counts but have the same number of MoE layers, leading to a similar absolute increase in fine-tuning time (211 vs. 226 seconds).
Table 4. Comparison of MoE and MoGE across different values of $k$ and $n$ in image classification.
Type $n=32$ $k=1$
$k\!=\!1$ $k\!=\!2$ $k\!=\!5$ $k\!=\!9$ $n\!=\!16$ $n\!=\!25$ $n\!=\!32$
MoE 55.38 53.46 53.58 52.07 55.51 55.22 55.38
MoGE 56.15 53.75 54.17 53.91 55.56 55.46 56.15
Dataset Model Type WCT TMC
Tiny-ImageNet ViT-S/8 MoE  40617  5295
MoGE  40865  5299
ViT-B/8 MoE 112826 12071
MoGE 113080 12074
SwinV1-T/7 MoE  17343  7163
ImageNet-1K MoGE  17554  7167
(fine-tuning) SwinV1-B/7 MoE  30791 14547
MoGE  31017 14549
@@ 4.4. Ablation Study
For our proposed MoGE, we set the hyperparameter $\lambda$ in Eq. (9) as described in Appendix B. We conduct the following ablation studies for the Gaussian lowpass filter.
Standard Deviation $\sigma$. We select the SwinV2-MoGE-T/4 model on CIFAR-100 from Table 2 under the scenario of training from scratch, with $n=32$ and top-1 routing. With the filter size $h$ fixed at 3, hyperparameter tuning is performed empirically through the following steps. First, the optimal $\sigma$ is identified by conducting a search while keeping $\sigma$ constant. Next, $\sigma_{0}$ is fixed at 10 and $\sigma_{\min}$ is set to the optimal $\sigma$ determined in the previous step. The search then focuses on identifying the optimal $\gamma$. Finally, using the optimal $\gamma$ obtained, $\sigma_{\min}$ is further refined through an additional search. In our experiments, the typical settings are $h=3$, $\sigma_{0}=10$, and $\gamma=0.3$. For each specific task, only $\sigma_{\min}$ requires tuning.
Table 6. Ablation study on the Gaussian lowpass filter.
@@
Filter Size $h$. We select the V-MoGE-S/8 model on Tiny-ImageNet from Table 2 under the scenario of training from scratch. Each MoGE layer consists of 128 experts with top-1 routing, and the corresponding 2D matrix is $Z\in\mathbb{R}^{8\times 16}$. For the average and Gaussian filters, we set $\lambda=4$e-3 and evaluate the performance across different values of filter size $h$. For the Gaussian filter, we further set $\sigma_{0}=10$, $\sigma_{\min}=1.5$, and $\gamma=0.3$. The results are summarized in the last two columns of Table 6, where “3a” and “5a” represent average filters with sizes $3\times 3$ and $5\times 5$, respectively, while “3g” and “5g” denote Gaussian filters with the same respective sizes. From the results, it is evident that filters with $h=3$ outperform those with $h=5$, and Gaussian filters consistently achieve better performance compared to average filters.
@@ 5. Conclusion
This paper introduces a novel perspective on MoE with the widely used top-$k$ routing, grounded in sparse representation. It proposes applying a group sparse regularization to the routing input, referred to as Mixture of Group Experts (MoGE), indirectly imposes structural constraints on the experts. MoGE further organizes the routing input into a 2D topographic map, grouping neighboring elements to encourage the learning of invariant representations. Consequently, MoGE fosters greater diversity and specialization among experts. Extensive experiments across various image classification and language modeling tasks demonstrate that the proposed MoGE consistently outperforms vanilla MoE, while incurring only minimal increases in computation and memory overhead. By scaling both the number of experts and the $k$ value in top-$k$ routing, MoGE offers a simple yet effective solution that enhances the flexibility of expert design in MoE models.
@@
Limitations. Our evaluation is currently limited to image classification and language modeling tasks. Due to constrained computational resources, we have not explored larger-scale MoE models with billions of parameters. Designing novel structural constraints in our MoGE framework tailored to specific tasks is a promising direction worth exploring. Additionally, the expert architecture in MoGE models is currently adapted directly from dense Transformer models, leaving the investigation of optimal expert design as an exciting direction for future research.

### v2

@@ 3.2. Language Modeling
Setup. Following (Teo and Nguyen, 2024), we evaluate the effectiveness of MoGE on WikiText-103 (Merity et al., 2017) using three types of MoE models: Switch Transformer (SMoE) (Fedus et al., 2022), MomentumSMoE (Teo and Nguyen, 2024), and AdamSMoE (Teo and Nguyen, 2024). MoGE introduces our proposed group sparse regularization to the MoE, while keeping the load balancing loss.
@@
Results. Table 3 summarizes the perplexity (PPL) values for three types of MoE models on the validation and the test sets of WikiText-103. Lower PPL values indicate better performance, with the best results highlighted in bold. It can be seen that both MoE and MoGE achieve better results with $k=2$ compared to $k=1$. MoGE consistently outperforms MoE across all test sets and most validation sets, clearly demonstrating its superiority. Although MoGE shows slightly inferior performance to MoE on the validation set when using the AdamSMoE-medium model, it still achieves better performance on the more critical test set. Due to space constraints, the training and the validation PPL curves are provided in Appendix A.
Table 3. Comparison of the perplexity (PPL) of MoGE across three types of MoE models.
@@ 3.3. Ablation Study
Scaling Expert Number $n$. We analyze the performance of MoGE when scaling $n$. For this investigation, we pretrain the SwinV2-MoE-T and SwinV2-MoGE-T models with 64 or 128 experts on Tiny-ImageNet and ImageNet-1K datasets, as in Table 7. The results consistently demonstrate that MoGE outperforms MoE. We also fine-tune SwinV1-MoE-B and SwinV1-MoGE-B with 128 experts on ImageNet-1K dataset, which comprise approximately 1.4B total parameters. As also shown in Table 7, MoGE achieves superior performance compared to MoE.
Efficiency. The additional computation and memory overhead of MoGE relative to MoE mainly depends on the expert number $n$, not the expert size or total model parameters. To validate this, we use four models from Table 2 and evaluate them in both pretraining and fine-tuning settings, using $n=32$ and top-1 routing. The results are summarized in Table 1 of Appendix A. The increase in GPU memory usage for MoGE over MoE is negligible. In terms of wall-clock time, pretraining ViT-based MoE models on Tiny-ImageNet with MoGE increases no more than 1%, while fine-tuning SwinV1-based models on ImageNet-1K increases no more than 1.5%. Despite large differences in model size (e.g., ViT-S vs. ViT-B or SwinV1-T vs. SwinV1-B), models with the same number of MoE layers exhibit similar absolute increases in training time (e.g., 248 vs. 254 seconds for ViT variants; 211 vs. 226 seconds for SwinV1 variants).
Table 4. Comparison of MoE and MoGE in scaling $n$.
Dataset Model $n$ $k$ Acc@1 Acc@5
Tiny-ImageNet SwinV2-MoE-T 128 1 58.14 78.12
SwinV2-MoGE-T 128 1 59.61 79.09
SwinV2-MoE-T 128 2 54.74 76.33
SwinV2-MoGE-T 128 2 59.12 78.91
ImageNet-1K SwinV2-MoE-T 64 1 80.07 94.97
SwinV2-MoGE-T 64 1 80.95 95.67
ImageNet-1K SwinV1-MoE-B 128 1 84.88 97.50
(fine-tuning) SwinV1-MoGE-B 128 1 85.75 97.98
Standard Deviation $\sigma$. We pretrain the SwinV2-MoGE-T model on CIFAR-100 (see Table 2), with $n=32$ and top-1 routing. With filter size $h=3$, hyperparameters are tuned as follows: we first search for the optimal $\sigma$, then fix $\sigma_{0}=10$ and set $\sigma_{\min}$ to this value before searching for the best $\gamma$. Using the optimal $\gamma$, we further refine $\sigma_{\min}$. In practice, we use $h=3$, $\sigma_{0}=10$, and $\gamma=0.3$, with only $\sigma_{\min}$ requiring tuning.
Filter Size $h$. We use the V-MoGE-S model on Tiny-ImageNet for pretraining. Each MoGE layer uses 128 experts with top-1 routing, yielding a $Z\in\mathbb{R}^{8\times 16}$ matrix. For both average and Gaussian filters, we set $\lambda=4$e-3 and evaluate different values of $h$. For Gaussian filters, $\sigma_{0}=10$, $\sigma_{\min}=1.5$, and $\gamma=0.3$ are used. Results are shown in the last two columns of Table 5, where “3a/5a” and “3g/5g” refer to $3\times 3$/$5\times 5$ average and Gaussian filters, respectively. Filters with $h=3$ outperform $h=5$, and Gaussian filters consistently surpass average filters.
Table 5. Ablations on the Gaussian lowpass filter.
@@ 4. Conclusion
This paper introduces a novel perspective on MoE with the widely used top-$k$ routing, grounded in sparse representation. It proposes applying a group sparse regularization to the routing input, referred to as Mixture of Group Experts (MoGE), indirectly imposes structural constraints on the experts. MoGE further organizes the routing input into a 2D topographic map, grouping neighboring elements to encourage the learning of invariant representations. Consequently, MoGE fosters greater diversity and specialization among experts. Extensive experiments across various image classification and language modeling tasks demonstrate that the proposed MoGE consistently outperforms vanilla MoE, while incurring only minimal increases in computation and memory overhead.
@@
Limitations. Our evaluation is currently limited to image classification and language modeling tasks. Designing novel structural constraints in our MoGE framework tailored to specific tasks is a promising direction worth exploring.
Appendix A Additional Results
A.1. Efficiency Analysis
Table 6 summarizes the computational and memory comparison between MoE and MoGE. The results show that MoGE introduces a minimal increase in wall-clock time and total GPU memory usage compared to MoE.
Dataset Model Type WCT TMC
Tiny-ImageNet ViT-S MoE  40617  5295
MoGE  40865  5299
ViT-B MoE 112826 12071
MoGE 113080 12074
SwinV1-T MoE  17343  7163
ImageNet-1K MoGE  17554  7167
(fine-tuning) SwinV1-B MoE  30791 14547
MoGE  31017 14549
A.2. Scaling $k$
We analyze the performance of MoE and MoGE when scaling $k$ in top-$k$ routing. For this investigation, we pretrain the V-MoE-S and V-MoGE-S models with 32 experts on the Tiny-ImageNet dataset. The results are summarized in Table 7. We can see that MoGE consistently outperforms MoE. Moreover, MoGE exhibits a less significant performance decline as $k$ increases. This highlights the ability of MoGE to leverage larger $k$ values.
Table 7. Comparison of MoE and MoGE across different values of $k$ on image classification.
Type $k\!=\!1$ $k\!=\!2$ $k\!=\!5$ $k\!=\!9$
MoE 55.38 53.46 53.58 52.07
MoGE 56.15 53.75 54.17 53.91
A.3. Scheduling Strategy for $\sigma$
As shown in Eq. (8), we adjust $\sigma$ to gradually decay from an initial value $\sigma_{0}$ to a minimum value $\sigma_{\min}$ as the iteration number $t$ progresses. The rate of this decay is governed by $\gamma$. Fig. 3 illustrates the effects of different $\gamma$ values on the decay behavior of $\sigma$, with $\sigma_{0}=10$, $\sigma_{\min}=2$, and the total number of iterations $T=45000$ fixed. Specifically, when $\gamma=0$, $\sigma$ remains constant throughout the training process. For $10<\gamma<$ 1, $\sigma$ decays rapidly in the early stages and then slows down as training progresses. In the case where $\gamma=1$, the decay is strictly linear across all iterations. Finally, for $\gamma>1$, the decay starts off slower but accelerates toward the later stages of training.
Figure 3. The curves of $\sigma$ with varying $\gamma$ over iterations.
Similar to the challenges of designing optimal learning rate scheduling strategies for neural networks, developing an effective $\sigma$ scheduling strategy is equally demanding, as it involves dynamic variations across multiple factors. Based on our empirical studies, it is a good choice to design $\sigma$ scheduling strategies that do not deviate significantly from $\gamma=1$. Future work could draw inspiration from learning rate scheduling techniques in neural networks to further refine $\sigma$ scheduling strategy.
A.4. Image Classification
Following previous works, we conduct testing using the model trained at the end of training on image classification. The training loss curves and top-1 accuracy curves corresponding to Table 2 of the main paper are presented in Figs. 4-11. Consistent with the conclusions in the main paper, MoGE incorporates a regularization constraint compared to MoE. Consequently, its training loss is not always lower than that of MoE. However, MoGE consistently achieves higher top-1 accuracy than MoE.
Figure 4. Training loss and top-1 accuracy curves of V-MoE and V-MoGE on CIFAR-100.
Figure 5. Training loss and top-1 accuracy curves of SwinV2-MoE and SwinV2-MoGE on CIFAR-100.
Figure 6. Training loss and top-1 accuracy curves of V-MoE and V-MoGE on Tiny-ImageNet.
Figure 7. Training loss and top-1 accuracy curves of SwinV2-MoE and SwinV2-MoGE on Tiny-ImageNet.
Figure 8. Training loss and top-1 accuracy curves of V-MoE and V-MoGE on ImageNet-1K.
Figure 9. Training loss and top-1 accuracy curves of SwinV2-MoE and SwinV2-MoGE on ImageNet-1K.
Figure 10. Training loss and top-1 accuracy curves of V-MoE and V-MoGE on ImageNet-1K (fine-tuning).
Figure 11. Training loss and top-1 accuracy curves of SwinV1-MoE and SwinV1-MoGE on ImageNet-1K (fine-tuning).
A.5. Language Modeling
Following previous studies (Teo and Nguyen, 2024), we conduct testing using the model that achieves the best performance on the validation set. It is worth noting that performance is evaluated during training only at the end of each epoch rather than after every iteration. Additionally, the test set is evaluated only once, resulting in a single test PPL value. As a result, we do not have test PPL curves. The training and validation perplexity (PPL) curves associated with Table 3 of the main paper are shown in Figs. 12-16. Although MoGE incorporates a regularization constraint compared to MoE, its training PPL remains either superior to or comparable to that of MoE. Likewise, its validation PPL is also either superior to or comparable to that of MoE.
Figure 12. Training and validation PPL curves of SMoE-small and SMoGE-small with $n=16$ and $k=1$.
Figure 13. Training and validation PPL curves of SMoE-small and SMoGE-small with $n=16$ and $k=2$.
Figure 14. Training and validation PPL curves of SMoE-medium and SMoGE-medium.
Figure 15. Training and validation PPL curves of MomentumSMoE-small (abbreviated as MomSMoE-small) and MomentumSMoGE-small (abbreviated as MomSMoGE-small).
Figure 16. Training and validation PPL curves of AdamSMoE-medium and AdamSMoGE-medium.
Table 8. Hyperparameters for image classification.
Settings CIFAR-100 Tiny-ImageNet ImageNet-1K ImageNet-1K (fine-tuning)
Batch size 128 128 1024 256
Optimizer AdamW AdamW AdamW AdamW
LR 1×$\times 10^{-3}$ 1×$\times 10^{-3}$ 5×$\times 10^{-4}$ 3×$\times 10^{-5}$
LR schedule cosine cosine cosine cosine
Weight decay 0.05 0.05 0.05 1×$\times 10^{-8}$
Warmup epochs 10 10 15 5
Epochs 150 150 150 30
Horizontal flip
Random crop
Auto augment
Mixup alpha 0.25 0.25 0.8 0.8
Cutmix alpha 0.25 0.25 1.0 1.0
Erasing prob. 0.25 0.25 0.25 0.25
Color jitter 0.12 0.15 0.4 0.4
Label smoothing 0.1 0.1 0.1 0.1
Gradient clip 5.0 5.0
Loss cross-entropy cross-entropy cross-entropy cross-entropy
Table 9. Training settings of MoGE for image classification.
Settings CIFAR-100 Tiny-ImageNet ImageNet-1K ImageNet-1K (fine-tuning)
V-MoGE SwinV2-MoGE V-MoGE SwinV2-MoGE V-MoGE SwinV2-MoGE V-MoGE SwinV1-MoGE
Resolution 32 32 64 64 224 192 224 224
Patch size 4 2 8 4 16 4 16 4
Embed 512 96 512 96 512 96 384 96
Blocks 8 [2,2,6,2] 8 [2,2,6,2] 8 [2,2,6,2] 12 [2,2,6,2]
Heads 8 [3,6,12,24] 8 [3,6,12,24] 8 [3,6,12,24] 6 [3,6,12,24]
Window - 4 - 7 - 12 - 7
Drop path rate 0.1 0.1 0.1 0.1 0.0 0.1 0.0 0.1
$n$ 32 32 32 32 32 32 32 32
$k$ 1 1 1 1 1 1 1 1
$\lambda$ 4×$\times 10^{-3}$ 4×$\times 10^{-3}$ 4×$\times 10^{-3}$ 4×$\times 10^{-3}$ $\times 10^{-5}$ 4×$\times 10^{-3}$ 4×$\times 10^{-7}$ 4×$\times 10^{-7}$
$\sigma_{0}$ 10 10 10 10 10 10 10 10
$\sigma_{\min}$ 1.5 1.5 1.5 1.5 0.5 0.5 0.5 0.5
$\gamma$ 0.3 0.3 0.3 0.3 0.3 0.3 0.3 0.3
Table 10. Training settings of MoGE for language modeling.
Settings Small Medium
SMoGE SMoGE MomentumSMoGE SMoGE MomentumSMoGE AdamSMoGE
#Layer 3 3 3 6 6 6
Hidden 128 128 128 352 352 352
Heads 8 8 8 8 8 8
Dropout 0.7 0.7 0.7 0.1 0.1 0.1
Batch size 96 96 96 48 48 48
Epochs 60 60 60 80 80 80
Optimizer Adam Adam Adam Adam Adam Adam
LR 1×$\times 10^{-3}$ 1×$\times 10^{-3}$ 1×$\times 10^{-3}$ 1×$\times 10^{-3}$ 1×$\times 10^{-3}$ 1×$\times 10^{-3}$
Warmup iterations 3000 3000 3000 4000 4000 4000
$n$ 16 16 16 16 16 16
$k$ 1 2 2 2 2 2
$\lambda$ 1×$\times 10^{-6}$ 1×$\times 10^{-6}$ 1×$\times 10^{-6}$ 1×$\times 10^{-6}$ 1×$\times 10^{-6}$ 1×$\times 10^{-6}$
$\sigma_{0}$ 10 10 10 10 10 10
$\sigma_{\min}$ 1.5 1.5 1.5 1.5 1.5 1.5
$\gamma$ 0.3 0.3 0.3 0.3 0.3 0.3
Appendix B Implementation Details
B.1. Datasets
Because only a single MoE layer is used, we utilize the Fashion-MNIST dataset (Xiao et al., 2017) to investigate how MoGE effectively learns invariant representations. For image classification tasks, we employ the CIFAR-100 (Krizhevsky and Hinton, 2009), Tiny-ImageNet (Le and Yang, 2015), and ImageNet-1K (Russakovsky et al., 2015) datasets. In language modeling tasks, the experiments are conducted using the WikiText-103 dataset (Merity et al., 2017).
The Fashion-MNIST dataset consists of 70,000 grayscale images, each with a resolution of $28\times 28$, representing fashion products across 10 categories, with 7,000 images per category. The dataset is divided into a training set of 60,000 images and a test set of 10,000 images.
The CIFAR-100 dataset comprises 60,000 color images, each with a resolution of $32\times 32$, distributed across 100 classes, with 600 images per class. Among these, 50,000 images are designated for training and 10,000 for testing, resulting in 500 training images and 100 testing images per class.
The Tiny-ImageNet dataset consists of 200 classes, each containing 500 training images, 50 validation images, and 50 test images. All images are resized to $64\times 64$ pixels. Following (Lee et al., 2021), we train our models from scratch using the training set.
The ImageNet-1K dataset comprises 1,000 object classes, including 1,281,167 training images, 50,000 validation images, and 100,000 test images. Following common practices (Liu et al., 2022a, 2021), most models are trained on the training set using a resolution of $224\times 224$. Follow (Hwang et al., 2023), a resolution of $192\times 192$ is adopted for SwinV2-based models (Liu et al., 2022a).
The WikiText-103 dataset is sourced from Wikipedia articles and is specifically designed to capture long-range contextual dependencies. The training set includes approximately 28,000 articles, amounting to a total of 103 million words, with each article segmented into text blocks of around 3,600 words. The validation and test sets consist of 60 articles each, containing 218,000 and 246,000 words, respectively, for a combined total of approximately 268,000 words.
B.2. Training Details for Exploring Invariant Representation Learning
[…]
During training, we use a batch size of 200 and optimize the models using the AdamW optimizer with a learning rate of 0.001. A cosine learning rate scheduler is employed, along with a weight decay of 0.05 and a warmup period of 10 epochs. The training process leverages cross-entropy as the loss function.
B.3. Training Details for Image Classification
Tables 8 and 9 present detailed information about the experimental setup for our MoGE method in image classification tasks. Specifically, Table 8 summarizes the hyperparameters used for CIFAR-100, Tiny-ImageNet, ImageNet-1K, and the fine-tuning process on ImageNet-1K. Meanwhile, Table 9 highlights the training settings for MoGE.
Additionally, we apply the load balancing losses proposed by (Riquelme et al., 2021), with the loss weight consistently set to 0.01. For all experiments, the random seed is fixed at 0. The AdamW optimizer is configured with epsilon = 1e-8 and $\text{betas}=(0.9,0.999)$. Furthermore, all model parameters are initialized using a truncated normal distribution within the range $[-2,2]$, with a mean of 00 and a standard deviation of $0.02$. These configurations are consistently applied to all the compared models.
B.4. Training Details for Language Modeling
For the language modeling tasks, we adopt the settings described in (Teo and Nguyen, 2024). Two model sizes are considered: small (3 layers) and medium (6 layers). The small models are trained for 60 epochs, while the medium models are trained for 80 epochs without incorporating the load balancing loss.
All model parameters are initialized using the Kaiming uniform distribution (He et al., 2015) with $a=\sqrt{5}$. Additional training details can be found in Table 10. Our reproduction of results and implementation of MoGE are based on publicly available source code at: https://github.com/rachtsy/MomentumSMoE.
B.5. Environments
Our experiments on the Fashion-MNIST, CIFAR-100, and Tiny-ImageNet datasets are conducted using a single NVIDIA A800-SXM GPU with PyTorch 1.13.1+CUDA11.8. For the ImageNet-1K dataset, we utilize four NVIDIA A800-SXM GPUs with PyTorch 1.13.1+CUDA11.8. Similarly, experiments on the WikiText-103 dataset are performed using four NVIDIA A800-SXM GPUs with PyTorch 2.0.1+CUDA11.8.